OpenStudy (anonymous):

A segment has endpoints at (-13, 19) and (4, -7). What is the x-coordinate of the midpoint of that segment? Enter your answer as a decimal. is there a formula that anyone knows for this kind of problem?

OpenStudy (anonymous):

(-13+4)/2 -4.5

OpenStudy (aroub):

(-7+19/2,4-13/2) =(5/2,-5/2) = ( 2.5,-2.5)

OpenStudy (anonymous):

thnx guys =)